Changes in Current Operating Assets and Liabilities—Indirect Method Huluduey Corporation's comparative balance sheet for current assets...

  1. Changes in Current Operating Assets and Liabilities—Indirect Method

    Huluduey Corporation's comparative balance sheet for current assets and liabilities was as follows:

    Dec. 31, Year 2 Dec. 31, Year 1
    Accounts receivable $18,000 $14,400
    Inventory 34,800 29,700
    Accounts payable 27,600 20,700
    Dividends payable 8,400 10,800

    Adjust net income of $160,000 for changes in operating assets and liabilities to arrive at net cash flow from operating activities.
    $

    Feedback

    Once you have calculated the changes in the current operating assets and liabilities determine what impact those changes would have on cash. For example if accounts receivable has increased from last year to this year does that mean the company has collected more cash or less cash? If accounts payable has decreased does that mean the company has more cash or less cash?

Check My Work

  • Previous
  • Next
Answer #1
Net income 160000
Adjustments to reconcile net income to
net cash flow from operating activities:
Changes in current operating assets and liabilities:
Increase in accounts receivable -3600 =14400-18000
Increase in inventory -5100 =29700-34800
Increase in accounts payable 6900 =27600-20700
Net cash flow from operating activities 158200